OpenStudy (anonymous):

the number of washers 3/32 inch thick that can be cut from a piece of stock 25.5 inches long, allowing 1/16 inch for waste for each cut is A. 160 B. 163 C. 260 D. 272 E. 408

OpenStudy (mertsj):



OpenStudy (mertsj):

What that picture illustrates is that the number of washers is 1 more than the number of cuts. Can you solve it now?

OpenStudy (anonymous):

i still dont get it

OpenStudy (mertsj):

If there are x cuts, there will be x+1 washers. So x cuts=1/6x

OpenStudy (mertsj):

x+1 washers = (3/32)(x+1)

OpenStudy (mertsj):

The total must be25.5 Write and solve the equation.